OpenStudy (anonymous):

If a man writes ‘QDGFHS’ for ‘TIGHER’, then in this way what will be write for ‘MONKEY’? (A) XDHNMK (B) YDJMLK (C) XDJMNL (D) XCHNMK pls tell me hoe to dese type of questions..

OpenStudy (anonymous):

guys pls help me...

Parth (parthkohli):

Aha!

Parth (parthkohli):

Q D G F H S | T I G H E R T is the letter after S. I is the letter after H. G is the letter after F H is the letter after G. E is the letter after D. R is the letter after Q.

Parth (parthkohli):

It works more like a mirror.

OpenStudy (unklerhaukus):

your good

OpenStudy (anonymous):

thnku sir...

Parth (parthkohli):

Monkey will be on the left.

OpenStudy (unklerhaukus):

x

Parth (parthkohli):

_ _ _ _ _ _ || M O N K E Y on the right**

OpenStudy (anonymous):

unkle- they hd given a best descp ..dnt u thnk sir?

Parth (parthkohli):

_ _ _ _ _ L || M O N K E Y a start will be this.

Parth (parthkohli):

You see how L is just before M? J And now just continue :)

OpenStudy (unklerhaukus):

pardon @gaurav_20?

Parth (parthkohli):

@UnkleRhaukus is better than me at Math :P

OpenStudy (unklerhaukus):

@ParthKohli is better than me at Math

OpenStudy (anonymous):

@UnkleRhaukus -ok..sir u all r best.....u all r gd...iam just giving clerification y ihd choosen best ans...thts it...we respect u all...

Parth (parthkohli):

I don't even know calculus... :I you're on every topic!

Parth (parthkohli):

@gaurav_20 Thank you!

OpenStudy (unklerhaukus):

yeah i made a few multiplication errors today

OpenStudy (anonymous):

hahaha.....sir its ok...if u do so sometimes...

Parth (parthkohli):

Everyone gets an error in multiplication problems like these\[23232824834734837434732 \cdot 8326838238416362363343139 \]

OpenStudy (anonymous):

@ParthKohli -nyc point sir....bt i say dnt solve ds...v wl get mad..

OpenStudy (unklerhaukus):

\[\sim1.9\times10^{47}\]

OpenStudy (unklerhaukus):

the error i made earlier today was \[-3\times-3=3?\]

Parth (parthkohli):

I see how you got that. Either counting digits in both numbers, or WolframAlpha/some damned software.

Parth (parthkohli):

I saw that one.... @UnkleRhaukus

OpenStudy (anonymous):

has anyone solved this problem yet?

Parth (parthkohli):

Yeah, @panlac01 it's done already

OpenStudy (anonymous):

I am too lazy to scroll through the response, what is the answer?

OpenStudy (anonymous):

responses*

OpenStudy (anonymous):

option c is correct

OpenStudy (anonymous):

reverse sequence after the actual letter of the word

OpenStudy (anonymous):

k

OpenStudy (anonymous):

@panlac01 u cn see the example in response

OpenStudy (anonymous):

I know how to answer. I was just wondering if it were the same as what others came up with.

OpenStudy (unklerhaukus):

\[23232824834734837434732 \cdot 8326838238416362363343139\] \[\sim(2.3\times10^{22})\times(8.3\times10^{24})\]\[=(2.3\times8.3)\times10^{22}\times10^{24}\]\[=[(2\times8.3)+3\times 0.1\times8.3]\times10^{22+24}\]\[=[(16.6)+3\times 0.83]\times10^{46}\]\[=[16.6+2.49]\times10^{46}\]\[=19.09\times10^{46}\]\[\approx1.9\times10^{47}\]

OpenStudy (anonymous):

ok....

Parth (parthkohli):

That... is... absolutely... brilliant...